Based on the following information below:

1. Calculate the direct materials price and quantity variance. Please note that the materials price variance is based on actual material purchased and the quantity variance is based on material used. Gourmet, Inc. prDduces containers of frozen food Duing October the company had the following actual production and costs Actu

2. Calculate the direct labor rate and efficiency variances.

3. Calculate the variable overhead spending and efficiency variances.

4. Calculate the fixed overhead budget variance.
Answer #1

Direct material price variance = Actual quantity purchased * (Standard price-Actual price)

Direct material price variance = 15,000 * ($2 - 28,000/15,000) = $1,950 Favorable

Direct material quantity variance = Standard price * (Standard quantity-Actual quantity)

Standard quantity = 720*20 = 14,400

Direct material quantity variance = $2 * (14,400-14,600) = $400 Unfavorable

Direct labor rate variance = Actual hours * (Standard rate-Actual rate)

Direct labor rate variance = 3,900 * ($19.75 - $80,000/3,900) = $2,964 Unfavorable

Direct labor efficiency variance = Standard rate * (Standard hours-Actual hours)

Standard hours = 720*5 = 3,600 hours

Direct labor efficiency variance = $19.75 * (3,600-3,900) = $5,925 Unfavorable

Variable overhead spending overhead = Actual hours * (Standard rate-Actual rate)

Variable overhead spending variance = 3,900 * ($1.50 - $5,700/3,900) = $156 Favorable

Variable overhead efficiency variance = $1.50 * (3,600-3,900) = $450 Unfavorable

Fixed overhead budget variance = Budgeted overhead - Actual overhead

Fixed overhead budget variance = $12,500 - 13,000 = $500 Unfavorable